OpenStudy (anonymous):

This matrix represents a system of equations. When fully row reduced, what will x equal? Assume the first column represents x. 1 3 -3 2 0 1 3 1 0 0 1 2 A. 23 B. -1 C. 11 C. 2

OpenStudy (anonymous):

Let the second column be y and the third column be z so the third row tells us that z=2 Then the second row says. y+3z=1 y+6=1 y=-5 And now to htre first row x+3y-3z=2 x -15-6 =2 x-21 =2 x=23